Home‎ > ‎Crew 164‎ > ‎

0226-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 5
Person filling out Report: Mohammad IRANMANESH, XO
Summary Title: Getting used to life on Mars. 
Mission Status: All systems GO. Blue ATV 2,3,4 and red one are OK. ATV 1 not tested because of axle issue signaled by Misison Support.

Sol Activity Summary: Morning sport routine. EVA. Started space suit gloves repair. Finished 3D design of water filter for upper deck with CATIA. NorCal rover succesfully teste during EVA. CRV test with new camera angle. Day with Phobos TV Journalist. Data recording on MOMA experiment was fianlly fixed.
Look Ahead Plan: First EVA with ATV tomorrow. Night at the observatory planned by Crew Astronomer. Concerning hab repairs: 3d printing of water filter, sewing gloves and building 4th backpack out of spares in progress. 

Anomalies in work: Internet data consumption anormally high today (probably due to Phobos TV Journalist smartphone connection to the HAB wifi network by mistake). We hope we will be able to go through the entire CAPCOM but not sure. We are sending reports early and will connect only one PC for CAPCOM just in case.
Small clean water running in front of the HAB. Probably due to a leak.
Weather: sunny, no clouds. Sky seems perfect for astronomy observations (except for moonlight).

EVA: 
Reports to be filed: 
- Commander report
- Engineering report
- Journalist report 
- EVA #5 report
- EVA #6 ATV request
- Astronomy report
Support Requested: 
- HAB Engineering Experts opinion on water leak observed (see Engineering Report).